OpenStudy (anonymous):

y = sin (e3x) the dy/dx =

OpenStudy (anonymous):

\[\sin(e^{3x})\]?

OpenStudy (anonymous):

yes

OpenStudy (ikram002p):

ok so remember this (f(g(x))'=f'(g(x)) g'(x) ok ?

OpenStudy (anonymous):

chain rule twice

OpenStudy (ikram002p):

yep

OpenStudy (akashdeepdeb):

\[y = \sin(e^{3x})\] \[\frac{dy}{dx} = \frac{d(\sin(e^{3x}))}{dx}\] Now use chain rule. \[\frac{d(\sin(e^{3x})) * d(e^{3x})}{d(e^{3x}) * dx}\] You'll get. [Use chain rule again in the next step] \[\cos(e^{3x})\frac{d(e^{3x})*d(3x)}{d(3x)*dx}\] Then you get: \[\cos(e^{3x}) ~*~ e^{3x} * 3\] Getting this? :)

OpenStudy (anonymous):

going step by step...

OpenStudy (akashdeepdeb):

See through the text. I have posted the steps. Are you ever of this rule called the chain rule?

OpenStudy (akashdeepdeb):

aware*

OpenStudy (akashdeepdeb):

No problem. I'll explain again.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the derivative of sine of something is cosine of something times the derivative of something the derivative of e to the power of something is e to the something times the derivative of something the derivative of a constant times x is the constant

OpenStudy (anonymous):

\[\frac{d}{dx}[\sin(\text {whatever})]=\cos(\text{whatever})\times \text{whatever}'\]

OpenStudy (akashdeepdeb):

Yes exactly! \[\frac{d}{dx}[cos(anything)] = ?\] What do we have here? We have to differentiate \(cos(anything)\) with dx. Is that possible? No. We can only differentiate \(cos(anything)\) with \(d(anything)\) not with dx. Because \[\frac{d}{d('x')}cos ('x') ~~gives~~us~~ = -sin('x')\] So why don't we multiply \(d(anything)\) to the numerator AND the denominator to make things right? So, we'll get: \[\frac{d}{dx}[cos(anything)] = -sin(anything) * \frac{d}{dx}(anything)\] [This is the chain rule] Understood? :)

OpenStudy (anonymous):

yes.. thank you.. :-)

OpenStudy (yanasidlinskiy):

By the way, \(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)